OpenStudy (anonymous):

Jack is selling holiday ornaments and boxes of candy. Each ornament is $1.50 and each box of candy is $3.00. Jack sold a total of 15 items for $27. How many boxes of candy did he sell?

OpenStudy (anonymous):

We can write a system of equations. 1.50x + 3.00y = 27 x + y = 15 Where x = ornaments, and y = boxes of candy.

OpenStudy (anonymous):

So can you solve that system of equations? @Kaylaaaaaaaaaa

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

how do i do that @iGreen

OpenStudy (anonymous):

Well we can use substitution. x + y = 15 Solve for x: Subtract y to both sides: x = -y + 15 Now we can plug in -y + 15 for x in the 1st equation: 1.50x + 3.00y = 27 1.50(-y + 15) + 3.00y = 27 Distribute 1.50 into the parenthesis: -1.50y + 22.5 + 3.00y = 27 Simplify: 1.5y + 22.5 = 27 Now subtract 22.5 to both sides, what's 27 - 22.5?

OpenStudy (anonymous):

Can you subtract that? @Kaylaaaaaaaaaa

OpenStudy (anonymous):

it is 4.5 @iGreen

OpenStudy (anonymous):

Yep, that gives us: \(1.5y = 4.5\) Now divide 1.5 to both sides, what's 4.5 / 1.5?

OpenStudy (anonymous):

3 @iGreen

OpenStudy (anonymous):

Yep, so we get: \(x = 3\) We said ornaments were 'x', so Jack sold 3 ornaments. It says Jack sold a total of 15 items. So to find out how many boxes of candy he sold, subtract 15 - 3, what do you get?

OpenStudy (anonymous):

12 @iGreen

OpenStudy (anonymous):

Yep, so Jack sold 3 ornaments, and 12 boxes of candies.

OpenStudy (anonymous):

thankyou @iGreen